# supplementary

### v2

@@ Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities between neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with widely disparate health outcomes. However, current statistical methods are often inadequate for appropriately defining what constitutes a spatial disparity and for constructing rankings of posterior probabilities that are robust under changes to such a definition. More specifically, non-parametric Bayesian approaches endow spatial effects with discrete probability distributions using Dirichlet processes, or generalizations thereof, and rely upon computationally intensive methods for inferring on weakly identified parameters. In this manuscript, we introduce a Bayesian linear regression framework to detect spatial health disparities. This enables us to exploit Bayesian conjugate posterior distributions in a more accessible manner and accelerate computation significantly over existing Bayesian non-parametric approaches. Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of mortality rates across tracheal, bronchus, and lung cancer.
@@ 1. Introduction
We propose methods for reporting spatial disparities on maps. Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023, and references therein) and spatial variation in health outcomes play a crucial role in detecting health disparities. While the role of spatial data analysis in epidemiological investigations has been extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and further references therein), inference pertaining to disparities comprises detecting “boundaries” on disease maps that identify significant differences among neighbors. These difference boundaries delineate regions with significantly different disease mortality and incidence rates and assists in improved geographic allocation of health care resources. In geographic analysis, the problem is often treated as one of “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches aiming to resolve this problem include, but are not limited to, developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra and gaoSpatialDifferenceBoundary2023.
In pursuing a probabilistic learning mechanism for adjacency boundaries on a map from data available as aggregates over regions, as is typical of health administration data, we recognize that administrative boundaries between adjacent regions need not represent health outcomes measured there. In somewhat contrarian manner, the more conspicuous disease mapping models assume that neighboring regions are similar to each other. This runs counter to detecting health disparities, where we seek out differences and not similarities. Therefore, modeling spatial disparities needs to balance underlying spatial similarities based upon geographic proximity with the ability to detect differences among neighbors. li2015bayesian, gaoSpatialDifferenceBoundary2023 and aiello2023detecting, for example, propose classes of models that endow the spatial random effects with a discrete probability law using Dirichlet processes. This allows for nonzero probabilities on spatial random effects being equal to each, which are computed as posterior probabilities and used to classify spatial difference boundaries by controlling a Bayesian false discovery rate (on the lines of mullerFDRBayesianMultiple2006).
We aim to obviate some challenges seen in the aforementioned approaches. For one, the computation of the Bayesian nonparametric models devised for difference boundary analysis is cumbersome with the presence of weakly identifiable parameters that impede convergence of iterative estimation algorithms. A second issue is that, while semi-parametric approaches offer some flexibility, the hierarchical Dirichlet process obscures the spatial dependence between the random effects and makes the inference sensitive to values of hyper-parameters. The analysis presented in gaoSpatialDifferenceBoundary2023 also reveal that such models may not deliver optimal sensitivity and specificity with regard to detection of boundaries. Furthermore, ascertaining difference boundaries using such models take into account only the probability of spatial effects being different from each other and not the differences between such effects. How small differences among neighboring effects translate to higher values of the posterior probabilities of the differences remain unaccounted. And the methods do not offer a defense against the argument that small differences in the values of the spatial effects of adjacent regions need not constitute a disparity.
We exploit conjugate distribution theory for hierarchical models to account for the differences between the values of the spatial effects as well as the posterior probabilities of such differences exceeding a certain threshold. In this regard, we extend the approaches in lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while accounting for false discoveries with the Bayesian FDR (mullerOptimalSampleSize2004).
@@
We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary and demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model. Section LABEL:sec:simulation_experiments evaluates our methods using simulations over a map of U.S. counties followed by an analysis of tracheal, bronchus and lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
2. Testing for Significant Contrasts in Bayesian Linear Regression
2.1. Conjugate Normal-Gamma Model
Consider the Bayesian linear regression model
$y=X\beta+\eta,\;\eta\sim N_{n}(0,\sigma^{2}V_{y})\;,$ (1)
where $y$ is an $n\times 1$ vector of observations, $X$ is an $n\times p$ matrix of predictors, $\beta$ is the unknown $p\times 1$ vector of corresponding regression coefficients, and $V_{y}$ is a $n\times n$ covariance matrix that is assumed to be completely specified by the analyst. A customary conjugate prior is
$\pi(\beta,\sigma^{2})=N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{0},b_{0})\;,$ (2)
where $M_{0}$ is a $p\times p$ prior correlation matrix for $\beta$ that is fixed, $m_{0}$ is a fixed $p\times 1$ vector so that $\mathbb{E}[\beta]=M_{0}m_{0}$, $a_{0}$ and $b_{0}$ are known shape and rate parameters of the inverse-Gamma prior on $\sigma^{2}$. Familiar Bayesian calculations yield the joint posterior distribution
$\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)\;,$ (3)
where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M=\left(M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X\right)^{-1}$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We study the problem of multiple pairwise comparison within the coefficients $\beta$ when $p>2$ and thus devote our attention to studying pairwise contrasts of the form $\beta_{i}-\beta_{j}$. The fundamental question is: after observing $y$, how can we make decisions about $\beta_{i}$ being different from $\beta_{j}$ for a large number of pairs $(i,j)$?
We must clarify how we intend to ascertain if $\beta_{i}$ is different from $\beta_{j}$. We are unable to simply use $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)$ since $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)=0$ for continuous models, such as (3), regardless of the observed data. Hence, it is not possible to directly evaluate the probability of $\beta_{i}=\beta_{j}$. One possibility is to use priors with point masses on the regression coefficients, such as slab-spike priors (mitchellBayesianVariableSelection1988) and several adaptations using mixture models to meaningfully evaluate $\mathbb{P}(\beta_{i}=0\,|\,y)$. For example, scottExplorationAspectsBayesian2006 investigate multiple comparisons of normal means to 00 by modeling each null hypothesis as independent Bernoulli trials with an unknown common failure probability. However, there is some difficulty in adopting a similar strategy for testing $\beta_{i}=\beta_{j}$ in the regression setting because posterior dependence among the elements of $\beta$ implies that it would be inappropriate to model the contrasts $\beta_{i}-\beta_{j}$ being zero as independent Bernoulli random variables. In other words, hypothesis testing of several contrasts on the regression coefficients is a multiple dependent hypothesis testing problem that prevents us from justifying a conditionally independent hierarchical model.
As it stands, the continuous conjugate prior under the standard Bayesian linear regression model is incompatible with difference probabilities that only consider exact equality. Changing the prior or the model sacrifices the analytical simplicity of the conjugate prior in the Bayesian linear regression setting, so we devote our attention to the third approach by considering a new class of difference probabilities. The goal here is to generate easily interpretable and informative statements about the parameters while also prioritizing model simplicity and computational cost.
@@ 2.2. Random Effects under an Augmented Linear System
The model in (1) can be criticized as overly simplistic if, for example, both spatial autocorrelation and unstructured random error substantially contribute to the true data generating mechanism. One model that attempts to address this is the BYM (Besag, York and Mollié) model, first introduced in besagBayesianImageRestoration1991, which includes both a spatially correlated and unstructured error component as part of a Poisson generalized linear model. However, simpsonPenalisingModelComponent2017 and rieblerIntuitiveBayesianSpatial2016 explain the error components are not identifiable and scaling issues complicate interpretation of the variance parameters in the BYM model. We follow the parametric form of the BYM model in rieblerIntuitiveBayesianSpatial2016 that ensures identifiable error components and imparts better interpretation. We write this BYM2 model as
@@
where $\phi$ is a $n\times 1$ vector of random effects, $\rho\in(0,1)$ controls the proportion of the total error attributed to the random effects’ variance structure, and $V_{\phi}$ is a known positive definite matrix. For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is a diagonal matrix with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for the parameters in (4), then it is helpful to express the joint posterior distribution in terms of the augmented model (banerjeeModelingMassiveSpatial2020a),
@@
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. An alternate representation that is interpreted in terms of least squares regression expresses $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ as the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and the scale parameter $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean of the residual sums of squares for the linear model in (5).
If $\rho\in(0,1)$ is fixed, then we can draw exact samples from the joint posterior. Denote $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$ and $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}$. Familiar calculations yield $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)\left(A^{-1}+A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}XA^{-1}\right)$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)/\rho\cdot A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y+(1-\rho)A^{-1}\left(X^{\mathrm{\scriptscriptstyle T}}B^{-1}X+I_{n}\right)m_{0}$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}})y-(1-\rho)B^{-1}XA^{-1}m_{0}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we have
@@
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}.$ (7)
A flat prior on $\beta$ considerably simplifies $\mathbb{E}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$ and $\operatorname{Var}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$. If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$l, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)\cdot(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y/\rho$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}\;.$ (8)
In (4), inference on differences within the fixed effects $\beta$ and the spatial residuals $\gamma$ are equivalent problems if $\rho$ is known. This implies that a decision framework for the Bayesian fixed effects model specified in (1) can be extended to draw inference on both fixed and random effects in the model given by (4) after conditioning on $\rho$ and $V_{\phi}$. Thus, we turn to the original problem of computing difference probabilities in the fixed effects model before extending to spatial disparities.

### v3

@@ Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities among neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes. However, there are statistical challenges to appropriately defining what constitutes a spatial disparity and to construct robust probabilistic inference for spatial disparities. We enrich the familiar Bayesian linear regression framework to introduce spatial autoregression and offer model-based detection of spatial disparities. We derive exploitable analytical tractability that considerably accelerates computation. Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
@@ 1. Introduction
We develop statistical methods for detecting spatial disparities on maps. Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps. Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein). Inference on disparities proceeds from “difference boundaries” that represent significant differences among neighbors and delineate regions with significantly different disease mortality and incidence rates, which assists in improved geographic allocation of health care resources. In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
While the aforementioned articles present an impressive range of methods, they are broadly classified into two groups. The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements. The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities. Both groups introduce spatial dependence and aim to balance underlying spatial similarities with the ability to detect differences among neighbors. Our current contribution offers a computationally scalable and analytically tractable approach to this problem. Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006). Benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in ascertaining spatial disparity, and delivering improved sensitivity and specificity for difference boundary detection (as seen in the analysis by gaoSpatialDifferenceBoundary2023).
@@
We exploit closed form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold. We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference. We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model. Section LABEL:sec:simulation_experiments evaluates our methods using simulated data over a map of U.S. counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
2. Assessing Contrasts in Bayesian Linear Regression
@@ 2.1. Random Effects under an Augmented Linear System
We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and afford better interpretation. We write this “BYM2” model as
@@
where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix. For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
@@
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
If $\rho\in(0,1)$ is fixed, then we can draw exact posterior samples from (3). Noting that $\gamma=\sigma\sqrt{\rho}\phi$ and carrying out some familiar linear algebra calculations, we obtain
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1},$ (4)
where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior on $\beta$, i.e., $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
@@
$\lim_{M_{0}^{-1}\to O}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-H\right)\right)^{-1},\;\mbox{ where }\;H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\;.$ (5)
These expressions are used in Section 3.2 to formulate difference boundary detection.